6x+ 4y = 20 000
2x+ 8y = 15 000

Answer by Edwin McCravy(20063)


First eliminate x and solve for y:

6x+ 4y = 20 000
2x+ 8y = 15 000

If we multiply the second equation through by -3
the -2x will become -6x and we can then add the
two equations term by term and the x-terms will 
cancel and be eliminated.  So we multiply 
2x+ 8y = 15 000 through by -3 and get 
-6x-16y = -45 000.  Then we write that
underneath the first equation, and add like terms
vertically and solve for y:

 6x+ 4y = 20 000
-6x-16y =-45 000
----------------
   -12y =-15 000
      y = (-15 000)/(-12)
      y = 1250

Now start over:

6x+ 4y = 20 000
2x+ 8y = 15 000

If we multiply the first equation through by -2
the 4y will become -8y and we can then add the
two equations term by term and the y-terms will 
cancel and be eliminated.  So we multiply 
6x+ 4y = 20 000 through by -2 and get 
-12x-8y = -40 000.  Then we write that
underneath the second equation, and add like terms
vertically and solve for x:

  2x+ 8y = 15 000
-12x- 8y =-40 000
-----------------
-10x     =-25 000
----------------
    -10x =-25 000
       x = (-25 000)/(-10)
       x = 2500

Edwin